the perimeter of a rectangle is 24. the width is 1 more than 1/4 it length. what is the length. THIS IS AN URGENT QUESTION!!!
Answer by addingup(3677)


2L+2W= 24
1/4L+1= W substitute in the first equation
2L+ 2(1/4L+1)= 24 Simplify and you get:
2L+ 0.5L+2= 24 Subtract 2, both sides:
2.5L= 22 Divide both sides by 2.5
L = 8.8